# Tailoring the Hyperparameters of a Wide-Kernel Convolutional Neural Network to Fit Different Bearing Fault Vibration Datasets

### v1

@@ 2 Background & Related Work
In this section, we review fault detection for the rotating machines seen in industrial settings. We also provide an overview of deep learning (DL) in the realm of time series analysis using sensor data, emphasising the design and training of convolutional neural networks (CNN). Additionally, given that this research concentrates on the architectural hyperparameters of a wide-kernel network, we discuss studies related to hyperparameter optimisation.
2.1 Fault Detection
@@
Detecting faults in rotating industrial machinery is essential to avoid breakdowns [19]. Fault detection data are typically collected from sensors that measure vibrations, represented as time series data. Initially, fault detection relied on physics-based models that required an understanding of the underlying mechanisms that generate and propagate vibrations [20]. However, these models struggled to adapt to changing environments and increasing data complexity. The advent of the Industrial Internet of Things (IIoT) and data-driven analysis techniques has revolutionised fault detection methods, enabling a more intelligent and automated approach [17, 21]. These advances eliminate the need for in-depth technical knowledge of industrial machinery, allowing for automated processing and adaptation to changing operational environments. For example, machine learning approaches such as K-NN [22, 23], Random Forest [24], and SVM [25, 26, 27, 28] have been applied for fault detection. However, these methods require extensive feature extraction, a time-consuming process that has to be fine-tuned towards the type of data used.
With the development of deep learning techniques, these steps were no longer necessary since deep learning techniques are able to automatically extract features from raw data, simplifying the otherwise complex feature extraction process. Initially, multilayer perceptrons (consisting of computational units arranged into simple layers where every unit in one layer connects to every unit in the next layer) were used, but their limited depth due to computational constraints shifted the focus to other architectures such as recurrent neural networks (RNNs), which can more effectively model temporal dependencies [29, 30]. An RNN processes sequences of data one step at a time, passing information along internally to keep track of what it has seen. The computational units in an RNN are arranged differently than for an MLP, since the RNN is formed from small blocks suited for processing sequences one step at a time whilst the layers of an MLP must process the entire input sequence in a single vast computation. However, RNN networks are less suited for time series data that appear as especially long sequences, which is common in sensor data sampled at a high frequency (e.g., mechanical vibrations or electrical currents). The main reason for this relates to the storage of long-term dependencies, also resulting in increased computational needs when the time series is of high resolution. On the other hand, convolutional neural networks (CNNs) including the one-dimensional version for processing time series gained popularity, demonstrating state-of-the-art performance in fault detection, especially with wide-kernel designs in the first convolutional layer [31, 32, 2, 33, 34]. Convolutions are an efficient alternative to RNNs because they also process sequences as a collection of smaller sub-pieces.
2.2 Deep Learning
Deep Learning draws inspiration from the human brain’s ability to combine many simple units into a large system capable of learning difficult tasks [35]. Its strength lies in processing and learning complex data, making it particularly effective for tasks requiring automatic feature extraction and refinement. This capability is crucial when valuable insights can be revealed by recognising patterns and relationships in the data  [36, 37].
The fundamental component of a neural network is the artificial neuron, or “node,” which processes multiple data inputs to produce an output. Many nodes can be grouped together into a layer and the term “deep” in deep learning refers to the inclusion of many layers in neural networks. A simple representation of a neuron $i$ is given by:
$a_{i}=\sigma\left(\sum_{j}w_{ij}x_{j}+b_{i}\right)$ (1)
Here, $x_{j}$ are the inputs, $w_{ij}$ are the weights, $b_{i}$ is the bias, and $\sigma$ is a nonlinear activation function. Each input is multiplied by a corresponding weight and the resulting values are added together along with an extra bias term. The formula operates somewhat like a recipe that combines different amounts of the various inputs (the $x_{j}$ terms) and adds some final garnish (the bias $b_{i}$). The nonlinearity introduced by the activation function is needed in order to handle difficult tasks, for example when similar inputs need different outputs. With layers stacking and interconnecting in complex architectures, neural networks can develop into intricate patterns, which gives them the flexibility and complexity needed to handle difficult tasks like image and speech recognition [36, 37, 38, 39].
The development of the multilayer perceptron (MLP) marked the beginning of deep learning, featuring fully connected layers [35, 40]. However, computational limitations initially restricted the depth of these networks. More advanced architectures were developed to overcome the limitations MLPs had when processing time series. Recurrent neural networks (RNNs), particularly long-short-term memory (LSTM) networks, emerged as effective architectures for time series analysis due to their ability to capture temporal dependencies [29, 30, 39]. Despite their success, RNNs are memory-intensive and less suited for long sequence data due to increased training times. RNNs are therefore not appropriate for directly processing raw inputs in many cases, and the undesirable step of pre-processing becomes necessary once again.
The initial application of MLPs in time series analysis spanned various domains, including stock prediction [41], weather forecasting [42], and fault detection [43]. Subsequently, RNNs [29, 39] and Convolutional Neural Networks (CNNs) [32, 21, 44, 17, 45] demonstrated significant performance improvements in time series tasks. CNNs, particularly one-dimensional (1D) CNNs were designed to process raw time series data effectively, leveraging automated feature extraction [31, 46]. These networks are robust against noise in the data and can be trained with relatively small datasets [47]. For fault detection a multi-layer perceptron (MLP) where all layers are fully connected [43] was initially proposed. Afterwards, the field of fault detection began to utilise convolutional neural networks (CNN) [17, 21, 44, 32] due to their notable performance improvements. CNN methods, when combined with data transformations like spectrograms, have been employed multiple times [48, 49]. However, 1D CNNs are able to process the raw time series data directly (without the need for spectrograms) and integrate automated feature extraction, and are often used in fault detection. These 1D CNNs also tend to be resilient to noise in time series data and can be trained with a small sample size [47].
@@ 2.2 Deep Learning
Subsequently, numerous refinements for fault detection using convolutions were introduced. These include, but are not limited to, wide-kernel CNNs [15, 33, 34] (which are used in this work), attention CNNs [50], periodic CNNs [51], and CNN autoencoders for unsupervised learning tasks [52]. Each of these networks has unique characteristics regarding complexity, computational requirements, and specific tasks. The general mechanics behind CNNs are further discussed in the next paragraph.
2.2.1 Convolutional Neural Networks
Convolutional Neural Networks (CNNs) are a specialised type of neural network that were originally created to handle two-dimensional (2D) data, such as images. Introduced by LeCun in 1989 [53, 38], CNNs utilise a feed-forward architecture that applies convolutions instead of general matrix multiplications, making them highly effective for data organised in a grid-like structure. In contrast to conventional MLPs, CNNs have become essential in areas such as Computer Vision, Image Recognition [38, 37], and are increasingly used in time series analysis.
@@ 2.2.1 Convolutional Neural Networks
The primary benefits of CNNs compared to other neural networks arise from their use of local receptive fields, weight-sharing, and sub-sampling [54]. These characteristics greatly decrease memory usage and computational complexity, thereby improving algorithmic efficiency. Convolutional layers apply filters to convolve input data and this basically allows the network to identify where patterns occur within a large signal. The activation function, commonly a Rectified Linear Unit (ReLU), introduces non-linearity, allowing the network to learn intricate patterns [55]. After convolution and activation, a pooling layer often shrinks the signal’s size in order to decrease the number of parameters of the next layer which helps prevent overfitting and lowers computational demands.
Initially, CNNs were less common for time series data due to their 2D nature, requiring conversion of one-dimensional (1D) data into 2D matrices, adding computational overhead. This changed with the development of 1D CNNs, which can directly process raw time series data [56, 34, 15, 47, 33, 31]. These advancements have solidified CNNs as a powerful tool for time series analysis, capable of efficiently handling high-frequency sensor data and extracting valuable features with minimal preprocessing.
@@
where $b_{i}^{l}$ is the bias, $k_{i}^{l}$ represents the filter weights, and $M^{l}(j)$ is the local region of the input in layer $l$. The weights of the filter act in some ways like a learned pattern, in the sense that the filter reacts most strongly (gives the largest numbers as outputs) when the input correlates with the weights of the filter in addition to being large. The filter applies to little regions of the overall signal with each region being one stride along from the previous one. In the notation above $i$ is a number used to refer to different regions. The most common activation function is ReLU [57], which introduces nonlinearity to the network.
Next, a pooling layer is utilised that allows the CNN to downsample the output of the convolutional layer making it more resource efficient in various applications from image and speech recognition to time series analysis.
2.3 Hyperparameter Search in Deep Learning
Past literature furnishes us with several examples of hyperparameter searches: activities in which authors tweak and tune the hyperparameters of a neural network in order to find the best performance. Because many decisions when designing and training neural networks can be called ‘hyperparameters’, previous studies have different degrees of overlap.
Working in the field of sentiment analysis, [58] used a grid search – which takes a list of valid values for each hyperparameter individually and then tests out every possible combination of these values – to optimise the accuracy of their LSTM-CNN network. Their motivation for using the grid search approach was to guarantee the absolute best performance, which more efficient search methods trade off in favour of increased processing speed. As will be described in Section 3, we also employ a grid search method, although our reason for doing so is that it supports a wider range of follow-on analyses rather than to find a single combination of hyperparameters that is best. Grid search is fairly common, for example being used by [59] to optimise the three training hyperparameters of learning rate, minibatch size and dropout ratio when training a CNN. The authors of [59] used the Talos package to implement their hyperparameter grid search.
Bayesian optimisation is another popular technique to find good hyperparameter settings. When developing a CNN-GRU network to detect different types of activities from wearable fitness tracker-type devices, [60] used the Optuna and OptKeras packages to implement their Bayesian optimisation search. The hyperparameters they tuned included the number of convolutional filters in different layers, and the number of hidden units in the GRU layers. The authors of [61] used SigOpt to perform Bayesian optimisation on a wide range of hyperparameters, from minibatch size and learning rate to the number of fully-connected neurons and the properties of the convolutional layers in their network. Furthermore, [62] applied Bayesian optimisation for LSTM-RNNs in highway traffic prediction.
Another popular possibility for performing a hyperparameter search is to use a genetic algorithm. At the beginning of a genetic algorithm, there is a population of candidate solutions which are tested out. The key step is that, after testing out a population of candidates, a new ‘generation’ of candidate solutions is created to replace the old population of candidates, and the new generation is formed by combining and mutating candidates from the previous generation with some preference being given to the candidates that performed best. Such an algorithm is used by [63] to tune hyperparameters such as the number of layers, the number and size of convolutions in a layer, and the number of fully-connected neurons in a CNN for identifying sleep apnea. A genetic algorithm was used by [64] to decide a very similar collection of hyperparameters, although in this case the CNN was applied to emotion recognition in smart tutoring.
Beyond the search methods already mentioned, there are more esoteric ideas that are sometimes presented as alternatives. An example of a more uncommon approach is to use orthogonal arrays to decide which hyperparameter combinations to test out. In [65], experiments with RNN and CNN networks on three time series datasets found that a hyperparameter search based on orthogonal arrays got closer to the global maximum level of accuracy for a limited time budget, compared to random search and Bayesian optimisation alternatives.
Although these studies provided valuable insights, their methods and objectives differ from the present work, which seeks a comprehensive view of network performance through a grid search rather than focusing on optimisation algorithms. Hence, our approach emphasises an extensive automated grid search, inspired by the work in [2, 1] to determine if the most important hyperparameters change as a result of differences in times series properties. In this way, we attempt to better understand and explain the hyperparameters wide-kernel CNN.
2.4 Summary of Own Previous Work
@@
In [2] we investigated how choices of the hyperparameters for a specific deep learning architecture (wide-kernel CNN) affect its performance in classifying faults from vibration signals in industrial machines. This architecture consists of five convolutional layers including an initial ‘wide’ layer with a large kernel size which then feed into a fully-connected classifier. See Section 3.3.1 for a detailed description. We were particularly interested in learning more about how performance changes across the hyperparameter space and so we performed a large-scale analysis, training and testing many versions of the wide-kernel CNN with different configurations (e.g., number of filters, kernel size) on three different datasets. The research used a grid search to explore the hyperparameter space and then presented multiple pieces of follow-on analysis.
@@
Looking at how the distribution of test accuracy scores changes as a result of tweaking individual hyperparameters, it was found that the number of filters in the later convolutional layers (3 to 5) has the biggest impact on performance, with an optimal value around 32. The kernel size in the first layer was also found to be important for datasets with a high sampling rate, where larger kernels perform better. Another discovery was that high-performing hyperparameter settings could be unstable, i.e., small changes to hyperparameters can have a large impact on performance, highlighting the need for careful tuning.
Overall, this study provided insights into how to configure a wide-kernel CNN for better fault detection in industrial machinery. We also identified interesting areas for future research, such as applying this approach to other signal classification tasks and different neural network architectures.
@@
A second study increased the number of benchmark datasets by four and applied 12,960 variations of the wide-kernel CNN to seven different industrial vibration datasets [1]. Once again, we analysed the grid search in order to identify how hyperparameter settings influence the network’s performance and how to effectively tune them for new datasets.
@@
3 Method
Our research method consists of several parts which feed into each other. In this section, we step through each piece of the method and describe it in detail. Starting out, we describe seven different benchmark datasets that we use to train and test neural networks. Different types of fault, machinery and recording set-up are possible in fault detection and we therefore expect that datasets can differ from one another. We want to account for variations in the data and so we assembled a selection of seven benchmarks to train and test out CNNs. The properties of the benchmark datasets are described in detail below.

### v2

@@ II Background & Related Work
In this section, we review fault detection for rotating machines in industrial settings. We also provide an overview of deep learning (DL) in the realm of time series analysis using sensor data, emphasising the design and training of convolutional neural networks (CNN). Additionally, given that this research concentrates on the architectural hyperparameters of a neural network, we discuss studies related to hyperparameter optimisation.
II-A Fault Detection
@@
Initially, fault detection relied on physics-based models that required an understanding of the underlying mechanisms that generate and propagate vibrations [13]. However, these models struggled to adapt to changing environments and increasing data complexity. The advent of the Industrial Internet of Things (IIoT) and data-driven analysis techniques has revolutionised fault detection methods, enabling a more intelligent and automated approach [11, 14]. These advances eliminate the need for in-depth technical knowledge of industrial machinery, allowing for automated processing and adaptation to changing operational environments. For example, machine learning approaches such as K-NN [15, 16], Random Forest [17], and SVM [18, 19, 20, 21] have been applied for fault detection. However, these methods require extensive feature extraction, a time-consuming process that has to be fine-tuned towards the type of data used. With the development of deep learning techniques, these steps were no longer necessary since deep learning techniques are able to automatically extract features from raw data, simplifying the otherwise complex feature extraction process.
Multilayer perceptrons (MLPs), also known as feedforward networks, are a relatively simple form of deep learning that have been applied to various domains of time series analysis, including stock prediction [22], weather forecasting [23], and fault detection [24]. Subsequently, recurrent neural networks (RNNs) [25, 26] and convolutional neural networks (CNNs) [27, 14, 28, 11, 29] demonstrated significant performance improvements in time series tasks.
@@
CNN methods, when combined with data transformations like spectrograms, have been employed multiple times [30, 31]. However, 1D CNNs are able to process the raw time series data directly (without the need for spectrograms) and integrate automated feature extraction, and are often used in fault detection [32, 33]. These 1D CNNs also tend to be resilient to noise in time series data and can be trained with a small sample size [34]. Over time, numerous refinements for fault detection using convolutions have been introduced. These include, but are not limited to, wide-kernel CNNs [35, 36, 37] (which are used in this work), attention CNNs [38], periodic CNNs [39], and CNN autoencoders for unsupervised learning tasks [40]. Each of these networks has unique characteristics regarding complexity, computational requirements, and specific tasks.
@@
The primary benefits of CNNs compared to other neural networks arise from their use of local receptive fields, weight-sharing, and sub-sampling [41]. These characteristics greatly decrease memory usage and computational complexity, thereby improving algorithmic efficiency. Convolutional layers apply filters to input data and this basically allows the network to identify where patterns occur within a large signal. The activation function, commonly a Rectified Linear Unit (ReLU), introduces non-linearity, allowing the network to learn intricate patterns [42]. After convolution and activation, a pooling layer often shrinks the signal’s size in order to decrease the number of parameters of the next layer which helps prevent overfitting and lowers computational demands.
@@
where $b_{i}^{l}$ is the bias, $k_{i}^{l}$ represents the filter weights, and $M^{l}(j)$ is the local region of the input in layer $l$. The weights of the filter act in some ways like a learned pattern, in the sense that the filter reacts most strongly (gives the largest numbers as outputs) when the input is large and correlates with the weights of the filter. The filter applies to little regions of the overall signal with each region being one stride along from the previous one. In the notation above $i$ is a number used to refer to different regions.
II-B Hyperparameter Search
Hyperparameter tuning appears most prominently in the field of ‘AutoML’, a research area based around the goal of finding and training the algorithm (which could be a neural network or another machine learning method) that will best perform a task, all without the involvement of humans. Although the hyperparameter space is considered one of the main preoccupations of AutoML researchers [43], it is often used simply to define which ranges of values will be considered ‘in scope’ when tuning an algorithm. However, better understanding the hyperparameter space gives us the chance to eventually build knowledge that generalises across many datasets instead of being restricted to concluding which algorithm is best on a narrow set of testing data [44]. The hyperparameter space is a way of specifying all the values the hyperparameters of an algorithm can take. If we combine this with extra information about how well the algorithm performs when using each combination of hyperparameters, we get the so-called “response landscape” or “response surface”.
For example, the response landscapes of neural networks are crucial to the discussion of random search by Bergstra and Bengio [45]. Their results suggest that many hyperparameters are not very important and thus it is not worth the time grid search (compared to random search) takes to try out each tuning option. This leads to the notion of “low effective dimensionality” in the response landscape, meaning that only a small number of changes need to be made to move from a low-performing set of hyperparameters to a high-performing set. With low effective dimensionality, random search is better than grid search. Additionally, the research concluded that different hyperparameters are important on different datasets. This implies dataset properties are an extra piece of the puzzle and the response landscape can change with the data. Bergstra and Bengio did not look further into dataset properties since doing so was not needed as part of their argument in favour of random search above grid search. However, we pick up on this line in our research, explicitly looking at the impact of the sampling rate and filtering of high frequencies within time series.
II-B1 Hyperparameter importance
Quantifying “hyperparameter importance” is another relevant sub-strand of past work which aims to probe the relationships between hyperparameters and neural network performance. Hutter, Hoos and Leyton-Brown [46] established what is probably the most widely-used method to estimate how important hyperparameters are. It uses a random forest as a ‘surrogate model’ predicting how well different hyperparameter combinations will perform. The surrogate model then leads to importance expressed as the variance explained (by the hyperparameter) within a functional ANOVA. The functional ANOVA has various nice properties including that it states the variance explained by subsets of hyperparameters as well as by individual hyperparameters. Moreover, the functional ANOVA makes it possible to ask ‘what performance is expected on average with Hyperparameter_A = value_X’. Subsequent research re-used the idea of using an fANOVA to find hyperparameter importance. For example, later work applied the technique to residual neural networks for image classification [44], and multi-objective tasks where the goal is to find a Pareto front balancing network performance against considerations like training time and energy consumption [47].
The other common way to assess hyperparameter importance is to perform ablation studies. In order to do an ablation study it is necessary to have already tuned the algorithm, since the goal is to find a path from the fully-tuned algorithm back to the basic default values for each hyperparameter. The path is generated by a greedy algorithm which, at each step, selects the hyperparameter which drops performance the most when it is reverted back to its default. The final path passes through all the hyperparameters in the order in which they contributed the most to performance [43]. Compared to the fANOVA approach, an ablation study does not include information about regions of the hyperparameter space that are distant from the defaults and the tuned values.
Compared to past work on hyperparameter importance, our own work has a number of differences worth pointing out, the first being that we focus on complex time series data analysed using modern neural network architectures. Time series datasets sophisticated enough to warrant neural networks are under-researched in connected to hyperparameter importance. We also use resampling and filtering to actively modify data rather than relying on pre-existing differences between datasets, because we place greater emphasis on understanding the context of bearing fault detection.
II-B2 Hyperparameter Search with Multiple Datasets
Finding knowledge that can generalise across datasets is one of the goals of AutoML and numerous approaches have been proposed for how to do it. Most commonly, the default values of an algorithm offer a way to transfer knowledge between datasets. Bayesian optimisation using surrogate models is a widely-used method which allows for transfer across datasets by ‘warm-starting’ the search process [9, 8]. Wistuba, Schilling and Schmidt-Thieme suggest a pruning method to leverage past experience from multiple datasets in order to speed up hyperparameter tuning [9]. Perrone et al. [48] trained a neural network to represent hyperparameter combinations in a way that makes predicting performance easier. The representations take hyperparameter values as inputs and transform them into 50-dimensional vectors that can be used as a basis for easily predicting how well the hyperparameters would do on a range of training datasets.
Considering the aforementioned sources, it is clear that generalising across datasets has been considered in past literature. However, it has taken the form of using black-box methods to find a single set of optimal values. By contrast, we are interested in getting to know the hyperparameters across a multiple values rather than identifying a single optimal point (and ignoring the rest of the hyperparameter space).
Much hyperparameter optimisation research has relied on Bayesian optimisation applied to surrogate models as a way to explore the hyperparameter space and find effective settings to use on new data. An interesting alternative comes in the form of “multiple defaults”, which basically means a small list of tried-and-tested combinations. These combinations are handed to a new user along with an ML algorithm, letting the user know what are sensible combinations of hyperparameter settings to try out. Arguments in favour of this approach include how easy to implement and use it is, and that the hyperparameter combinations can be tried in parallel. When trying to leverage knowledge from multiple datasets, this approach has the nice property that it is possible to make sure that at least one of the default settings performs well on each benchmark dataset. Multiple defaults is a concept suggested by Wistuba et al. [49], later developed into a method based on greedy forward search by Pfirsterer et al. [8], who chose 6 algorithms from OpenML to test out their method on 100 datasets of between 500 and 100000 observations and up to 5000 features. Both papers did not use any neural network methods however, unlike our research which focuses specifically on a wide-kernel convolutional architecture.
II-B3 Neural Architecture Search
Architectural hyperparameters make it possible to modify the shape and structure of a neural network before using it. Within AutoML, the niche that focuses on architectural hyperparameters is known as “neural architecture search” (NAS) [43] and it is often decomposed further into two types of decision: (1) choice of an overall structure and (2) choice of sub-modules. These are called macro and micro searches. Micro searches can rely on a template for the full network, like ResNet which is highly repetitive, in order to take a module and convert it into a testable network. Avoiding the combinatorial problem of looking at all the ways different layers can combine, micro searches can potentially offer a smaller search space than macro searches are liable to. To pick just one example of many, NASNET is a micro search. This research considered options for ‘normal’ and ‘reduction’ components in convolutional networks in order to identify damage caused by COVID-19 in chest x-ray images, and ended up achieving 97% accuracy [50]. Our work focuses on architectural hyperparameters within a relatively fixed network architecture, reflecting adaptations of the network rather than the design of a completely new network. This allows us to keep within a restricted search space and focus on understanding hyperparameters, a crucial consideration that is generally overlooked in the neural architecture search literature. Another point to note is that our work focuses on time series data, which has been much less investigated than other types of data in the context of NAS [50].
II-B4 Meta-Learning
Amongst past work on hyperparameter optimisation, there is one more trend that is relevant to our current study. This trend is called “meta-learning” and its goal is to predict which ML algorithm will perform a new task well, using a ‘meta’ algorithm to make the decision. The key information used to make predictions is the observed performance of ML algorithms on past datasets along with “meta-features” which summarise (previous and new) datasets. Meta-features describe the properties of a dataset and can be used to say how similar two datasets are. A trained “meta-model” can be used to either directly suggest an ML algorithm to try out on a new dataset, or it can be used to warm-start a search method such as Bayesian optimisation [43].
Different types of meta-feature exist, ranging from simple metadata like the number of variables measured, to parameters extracted from fitted statistical models, or even the performance scores of quick-to-evaluate models like linear regression [51]. One of the main goals of our research is to investigate how time series data properties relate to architectural hyperparameters. Probably the closest past research to our own is [52], which used meta-learning to predict the performance of algorithms in a time series prediction challenge. They extracted features specific to time series using the TSFresh and catch22 analysis packages. Similarly, in order to predict COVID-19 instances across different countries, [53] also used time series features for meta-learning although that research focused more on simpler ML algorithms.
A key distinction of our work is that we focus on two of the most understandable meta-features (sampling rate and presence of high frequencies) and we directly manipulate these within our datasets. Previous research into meta learning has relied on coincidental differences in existing datasets, without actively manipulating meta features like the sampling rate. Additionally, our research focuses on the domain of bearing fault detection in vibration data, within which the differences between datasets might be more subtle than they are in more generic meta-learning benchmarks, while at the same time presenting larger and more challenging datasets than most time series benchmarks for meta-learning.
II-C Summary of Own Previous Work
@@
In [2] we investigated how choices of the hyperparameters for a specific deep learning architecture (wide-kernel CNN) affect its performance in classifying faults from vibration signals in industrial machines. This architecture consists of five convolutional layers including an initial ‘wide’ layer with a large kernel size which then feed into a fully-connected classifier. See Section III-C1 for a detailed description. We were particularly interested in learning more about how performance changes across the hyperparameter space and so we performed a large-scale analysis, training and testing many versions of the wide-kernel CNN with different configurations (e.g., number of filters, kernel size) on three different datasets. The research used a grid search to explore the hyperparameter space and then presented multiple pieces of follow-on analysis.
@@
Looking at how the distribution of test accuracy scores changes as a result of tweaking individual hyperparameters, we found that the number of filters in the later convolutional layers (3 to 5) has the biggest impact on performance, with an optimal value around 32. The kernel size in the first layer also appeared to be important for datasets with a high sampling rate, where larger kernels perform better. Another discovery was that high-performing hyperparameter settings could be unstable, i.e., small changes to hyperparameters can have a large impact on performance, highlighting the need for careful tuning.
@@
A second study increased the number of benchmark datasets by four and applied 12,960 variations of the wide-kernel CNN to seven different industrial vibration datasets [1]. Once again, we analysed the grid search to identify how hyperparameter settings influence the network’s performance and how to effectively tune them for new data.
@@
Compared to our past work, the current study is able to make a broader argument about the importance of hyperparameter tuning by considering two alternative architectures to a wide-kernel CNN. Namely, we now experiment with a long- short-term memory (LSTM) and a transformer architecture. This allows us to show that hyperparameters are important across numerous fault detection algorithms, and also that hyperparameter tuning is important when neural networks are used across seven different fault datasets.
In addition, we investigate the impact of actively modifying dataset properties via resampling and filtering. This leads to new experiments with modified versions of one benchmark dataset. Ultimately our experiments with resampling and filtering lead to the new conclusion that sampling rate and the presence of high-frequency noise are not the main reason that a wide kernel is better than a short kernel, contrary to earlier speculation.
Moreover, we also conclude that due to the impact of differences between datasets, there will be no single set of default hyperparameter settings that is always best. We therefore establish “multiple defaults” for our wide-kernel CNN architecture, an extremely efficient way to try out hyperparameters combinations on new data which we did not explore in our previous research.
III Method
Our research method consists of several parts which feed into each other. In this section, we step through each piece of the method and describe it in detail. Starting out, we explain that we use seven different benchmark datasets in order to account for variations in the data used to train and test neural networks.